Considere the ideal I=<x²-4x+3,x³+3x²-x-3> of the ring Q[x]. Find a polynomial p such that I=<p>.Is Q[x]/I a field? Give reasons for your answer.
1
Expert's answer
2020-10-20T18:51:00-0400

Since "\mathbb Q[x]" is a Principal ideal domain, "p" is the gratest common divisor of the polinomials "x^2-4x+3" and "x^3+3x^2-x-3". Taking into account that "x^2-4x+3=x^2-x-3x+3=x(x-1)-3(x-1)=(x-3)(x-1)" and

"x^3+3x^2-x-3=x(x^2-1)+3(x^2-1)=(x+3)(x^2-1)=(x+3)(x+1)(x-1)", we conclude that "p=x-1".


Consider the ring "\mathbb Q[x]/I=\mathbb Q[x]/\langle x-1\rangle".


Polynomial remainder theorem implies that each element of "\mathbb Q[x]/\langle x-1\rangle" is of the form "[f(x)]=f(x)+\langle x-1\rangle=f(x)+(x-1)\mathbb Q[x]", where "\deg f(x)<\deg(x-1)=1", and therefore, "f(x)\in\mathbb Q". Consequently, "\mathbb Q[x]/\langle x-1\rangle=\{[a]\ : \ a\in\mathbb Q\}".


Consider the map "\psi: \mathbb Q\to \mathbb Q[x]/\langle x-1\rangle, \psi(a)=[a]." Taking into account that "\psi(ab)=[ab]=[a][b]=\psi(a)\psi(b)" and "\psi(a+b)=[a+b]=[a]+[b]=\psi(a)+\psi(b)", we conclude that "\psi" is a homomorphism. If "a\ne b", then "\psi(a)=[a]\ne [b]=\psi(b)", and consequently "\psi" is injective. Since "\psi(a)=[a]" for any "[a]\in \mathbb Q[x]/\langle x-1\rangle", "\psi" is surjective. Therefore, "\psi" is a isomorphism. Thus, "\mathbb Q" and "\mathbb Q[x]/\langle x-1\rangle" are isomorphic.


Since "\mathbb Q" is a field, we conclude that "\mathbb Q[x]/I" is a field.